# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ 3.3 FedP2EFT: FL to Personalize PEFT
The limited data available to each client in FL makes it difficult to train an effective PS in isolation, frequently resulting in overfitting. Following FedL2P (Lee et al., 2023), we mitigate this by federatedly learning a common PSG that generates client-wise PS. Concretely, we use a small, one hidden layer multilayer perceptron (MLP) with parameters $\phi$ that takes as input the client meta-data and outputs an estimated PS as follows:
@@
Federated Training of FedP2EFT. Fig. 2 shows the entire FedP2EFT algorithm during federated training. For each federated round, each sampled participating client $i$ receives $\phi$ from the server and loads them into its MLP. They then 1 perform a forward pass of the local train dataset on their base model and a forward pass of the MLP with the resulting features as per Eq. (3.3). 2 The estimated $\bm{\hat{\lambda}}^{i}$ is plugged into our proposed BayesTune-LoRA (Section 3.2) and 3 fine-tuning is performed as per Eq. (1) for $s$ steps (Stage 1). 4 The resulting $\hat{\bm{\lambda}}^{i,s}$ is used as an approximated ground-truth for regression learning of MLP to target the learned $\hat{\bm{\lambda}}^{i,s}$, where $\mathcal{L}_{1}$ is the L1 loss (Stage 2). Finally, 5 $\phi$ is sent back to the server for aggregation. As there is no single aggregation method that outperforms all others in every situation (Matsuda et al., 2022; Chen et al., 2022; Ye et al., 2024), we utilize FedAvg (McMahan et al., 2017). The aggregated $\phi$ is then sent to clients for the next round.
At the end of federated training, the learned $\phi$ can be deployed to any client, seen or unseen during federated training. Note that unlike FedL2P, which requires federated training for every target rank, FedP2EFT inherits the property of BayesTune-LoRA; federated training is a one-time cost for all ranks $\leq r_{\text{max target}}$.
Inference with FedP2EFT. Fig. 1(a) shows how FedP2EFT personalizes PEFT for each client upon deployment. Given the learned MLP and the client’s base model, 1 the client meta-data are retrieved (Eq. 3.3) and used to generate the client’s PS, $\bm{\lambda}$. 2 Given the client’s resource budget of total rank $r\cdot L$, we take the top-$(r\cdot L)$ largest ranks in $\bm{\lambda}$, freeze them, and initialize our proposed BayesTune-LoRA modules for all layers where $\|\lambda_{l}\|_{1}$$>$00. 3 The personalized LoRA ranks are used for fine-tuning before merging back to the base model to obtain the final personalized model.
Figure 2: FedP2EFT’s federated training of PSG for each federated round (See text in Section. 3.3).
@@ 4 Evaluation
4.1 Experimental Setup
We conduct experiments on multilingual scenarios, where clients with diverse high- and low-resource languages can collaboratively learn how to personalize a given base model to better cater to their language preferences. In all experiments, we divide clients in two pools, seen and unseen, where only the clients in the seen pool actively participate in federated training. We set the maximum number of communication rounds for training the PSG to $150$, randomly sampling $10\%$ of participating clients every round. We use Adam (Kingma & Ba, 2015) as the default optimizer for all our experiments. We evaluate on resource budgets $r=2,4,8,16$ where the total rank budget is $r\cdot L$. We summarize the FL scenarios that we consider in our experiments, leaving comprehensive details in Appendix A.
4.1.1 Tasks, Models, and Datasets
Text Classification. We adopt the pretrained multilingual BERT (Devlin et al., 2018) (mBERT) for all text classification experiments. For datasets, we introduce additional data heterogeneity to the simulated FL setups, XNLI (Conneau et al., 2018) and MasakhaNEWS (Adelani et al., 2023), proposed in PE_FL (Zhao et al., 2024).
For our XNLI setup, we sample 2k instances for train and $500$ for test in each pool. In contrast to PE_FL, which had $15$ clients ($1$ language per client), we divide the data equally among $20$ clients for each language. We then adopt the latent Dirichlet allocation (LDA) partition method (Hsu et al., 2019; Yurochkin et al., 2019), $y\sim Dir(\alpha)$, to simulate non-IID label shifts among these clients, with $\alpha=0.5$. Hence, there is a total of $600$ clients ($15$ languages $\cdot$ $20$ clients per language $\cdot$ $2$ pools), consisting of both label and feature heterogeneity.
@@ 4.1.1 Tasks, Models, and Datasets
Instruction-Tuning Generation. We adopt pretrained MobileLLaMA-1.4B (Chu et al., 2023) and Llama-3.2-3B (Dubey et al., 2024), which are representative of commonly supported model sizes on recent high-end edge devices (Mehta et al., 2024; Tan et al., 2024; Xu et al., 2024a). For each model, we run experiments on the recent Fed-Aya dataset. Fed-Aya is a real-world FL dataset naturally partitioned by annotator ID and each client has data with up to $4$ languages. Out of a total of $38$ clients, we select $8$ clients for our unseen pool. We also split each client’s data $80\%/20\%$ for train and test, respectively. Fig. 3 shows the distribution of predominant languages, where predominant refers to the client’s most commonly used language, in our setup.
Figure 3: The number of clients in each predominant language in our Fed-Aya setup.
4.1.2 Complementary Approaches
We show FedP2EFT’s compatibility with both off-the-shelf models and models trained using existing FL methods. Concretely, given a pretrained model, we obtain a base model using one of the following approaches:
Standard FL. We further train the pretrained model federatedly on the seen pool, either using existing PEFT methods or full fine-tuning (Ye et al., 2024; Sun et al., 2024),
Personalized FL. We adopt two recent personalized FL works: i) FedDPA-T (Yang et al., 2024b), which learns per-client personalized LoRA modules in addition to global LoRA modules, and ii) DEPT (SPEC) (Iacob et al., 2025), which learns per-client personalized token and positional embeddings while keeping the rest of the model shared. The base model hence differs for each client.
@@ 4.1.2 Complementary Approaches
Table 1: Mean±SD Accuracy of each language across 3 different seeds for seen clients of our MasakhaNEWS setup. The pretrained model is trained using Standard FL with full fine-tuning and the resulting base model is personalized to each client given a baseline approach.
$\mathbf{r}$ Approach eng som run fra lin ibo amh hau pcm swa orm xho yor sna lug tir Wins
2 LoRA 90.44±0.10 60.09±0.32 81.37±0.51 88.63±0.00 83.53±0.54 79.83±0.24 45.74±0.00 75.79±0.00 96.05±0.00 78.99±0.00 64.20±0.00 69.14±0.32 79.18±0.23 78.80±0.00 67.57±0.00 44.85±0.00 0
AdaLoRA 89.87±0.00 59.41±0.32 81.16±0.29 88.63±0.00 82.76±0.00 78.97±0.00 45.21±0.00 75.05±0.15 96.05±0.00 78.57±0.00 63.99±0.29 69.59±0.00 78.86±0.23 78.80±0.00 67.57±0.00 44.85±0.00 0
BayesTune-LoRA 89.87±0.00 59.63±0.32 81.37±0.00 88.63±0.00 83.14±0.54 78.97±0.00 45.21±0.00 75.16±0.00 96.05±0.00 78.57±0.00 63.99±0.29 69.59±0.00 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
FedL2P 90.72±0.59 61.00±1.16 81.99±0.88 89.10±0.67 83.91±0.00 79.66±0.24 45.74±0.00 76.73±1.12 96.05±0.00 79.69±0.40 64.40±0.29 69.14±0.32 79.67±0.61 78.80±0.00 67.87±0.42 45.34±0.69 0
FedP2EFT 91.98±0.00 65.54±1.16 87.79±0.29 93.52±0.23 88.51±0.94 82.56±0.00 51.95±0.25 79.66±0.29 97.59±0.31 84.73±0.52 72.22±1.01 76.35±0.55 82.11±0.23 81.34±0.68 69.07±0.42 63.48±0.34 16
4 LoRA 91.35±0.00 61.90±0.00 83.23±0.00 89.10±0.00 84.68±0.54 80.17±0.24 47.34±0.00 77.57±0.15 96.05±0.00 79.83±0.34 64.61±0.29 69.37±0.32 80.49±0.00 78.80±0.00 67.87±0.42 45.83±0.34 0
AdaLoRA 89.87±0.00 59.63±0.32 80.96±0.29 88.63±0.00 83.53±0.54 78.97±0.00 45.21±0.00 75.47±0.00 96.05±0.00 78.57±0.00 63.79±0.29 69.59±0.00 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
BayesTune-LoRA 89.87±0.00 59.41±0.32 80.96±0.29 88.63±0.00 82.76±0.00 78.97±0.00 45.21±0.00 75.47±0.26 96.05±0.00 78.71±0.20 64.20±0.00 69.59±0.00 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
FedL2P 91.70±0.40 63.49±1.16 83.23±0.51 90.84±0.97 83.91±0.94 80.00±0.42 47.69±1.40 79.45±1.07 96.05±0.00 81.51±1.03 64.81±0.50 71.62±1.99 81.46±0.80 78.80±0.00 68.17±0.85 46.32±0.60 0
FedP2EFT 92.05±0.44 67.12±0.64 86.34±0.00 93.36±0.39 91.19±1.43 82.56±0.84 53.90±0.25 81.66±0.65 98.46±0.31 83.89±1.10 72.84±0.51 77.70±0.00 82.11±0.23 83.70±0.00 74.47±0.42 67.65±0.00 16
8 LoRA 91.56±0.00 63.95±0.00 82.82±0.29 90.84±0.23 86.21±0.00 79.83±0.24 49.82±0.25 78.83±0.15 96.05±0.00 81.79±0.52 65.64±0.29 72.30±0.55 81.79±0.23 79.89±0.00 67.87±0.42 44.85±0.00 0
AdaLoRA 89.87±0.00 59.86±0.00 80.96±0.29 88.63±0.00 83.14±0.54 78.97±0.00 45.21±0.00 75.37±0.15 96.05±0.00 78.85±0.20 64.20±0.00 69.59±0.00 78.86±0.23 78.80±0.00 67.57±0.00 44.85±0.00 0
BayesTune-LoRA 89.94±0.10 59.41±0.32 80.75±0.00 88.63±0.00 83.14±0.54 78.97±0.00 45.21±0.00 75.47±0.26 96.05±0.00 78.71±0.20 64.20±0.00 69.59±0.00 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
FedL2P 91.91±0.26 64.18±0.32 83.02±0.58 91.15±0.59 84.29±0.54 80.51±0.00 48.58±0.67 79.98±0.82 96.27±0.31 82.49±0.79 66.25±1.62 72.07±1.39 81.95±1.05 79.16±0.51 68.47±1.27 47.79±3.65 1
FedP2EFT 91.42±0.20 68.03±0.96 85.09±0.00 92.42±0.39 91.95±0.00 81.88±0.24 55.32±0.43 82.70±0.26 98.68±0.00 84.31±0.20 73.25±0.58 78.15±0.32 84.72±0.46 85.15±0.26 74.17±0.42 68.38±0.60 15
16 LoRA 91.91±0.10 64.63±0.00 84.47±0.00 91.47±0.00 86.21±0.00 81.54±0.00 55.85±0.43 81.87±0.15 96.27±0.31 84.45±0.00 73.66±0.77 73.20±0.32 82.60±0.61 80.80±0.26 70.87±0.42 52.94±0.00 3
AdaLoRA 89.87±0.00 59.86±0.00 80.96±0.29 88.63±0.00 82.76±0.00 78.97±0.00 45.21±0.00 75.05±0.39 96.05±0.00 78.85±0.20 63.79±0.29 69.59±0.00 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
BayesTune-LoRA 90.08±0.18 59.63±0.32 81.16±0.29 88.63±0.00 83.53±0.54 78.97±0.00 45.21±0.00 75.58±0.15 96.05±0.00 78.57±0.00 64.20±0.00 69.37±0.32 79.02±0.00 78.80±0.00 67.57±0.00 44.85±0.00 0
FedL2P 91.70±0.10 66.22±0.32 85.09±0.51 92.10±0.23 86.21±0.00 80.34±0.24 54.79±1.15 81.13±0.26 96.71±0.00 83.89±0.40 71.81±2.04 75.68±0.00 84.06±0.23 80.61±0.68 70.87±0.85 57.84±4.86 2
FedP2EFT 91.14±0.30 67.80±0.85 84.47±0.00 91.31±0.22 91.95±0.00 81.54±0.42 51.95±0.25 83.44±0.39 98.25±0.31 82.49±0.71 76.54±0.50 79.50±0.32 84.39±0.00 85.87±0.00 75.38±0.43 68.38±0.00 10
$\mathbf{r}$ Approach eng som run fra lin ibo amh hau pcm swa orm xho yor sna lug tir Wins
2 LoRA 90.72±0.00 68.48±0.32 81.99±0.51 84.36±0.00 79.55±0.00 76.75±0.24 46.28±0.00 75.45±0.15 90.20±0.00 75.63±0.34 61.96±0.00 64.21±0.63 80.10±0.00 74.59±0.00 65.18±0.00 41.91±0.00 1
AdaLoRA 90.30±0.00 67.35±0.00 80.75±0.00 84.36±0.00 79.55±0.00 76.92±0.00 46.28±0.00 74.82±0.15 90.20±0.00 75.91±0.20 61.55±0.29 64.21±0.32 79.61±0.00 74.59±0.00 65.18±0.00 41.91±0.00 0
BayesTune-LoRA 90.30±0.00 67.35±0.00 80.75±0.00 84.36±0.00 79.55±0.00 76.92±0.00 45.92±0.25 74.61±0.00 90.20±0.00 75.49±0.20 61.55±0.29 63.76±0.00 79.45±0.23 74.41±0.25 65.18±0.00 42.65±0.00 0
FedL2P 90.58±0.10 68.48±0.32 82.61±0.00 84.36±0.00 79.55±0.00 76.41±0.42 46.81±0.75 75.97±0.59 90.20±0.00 75.91±0.40 62.17±0.29 64.43±0.95 80.26±0.23 74.77±0.26 66.07±0.00 41.91±0.00 0
FedP2EFT 90.65±0.10 72.11±0.56 88.41±0.29 88.15±0.39 86.36±0.00 81.54±0.42 51.06±0.00 79.52±0.30 93.68±0.31 78.01±0.40 73.01±0.50 78.52±1.10 83.66±0.23 80.00±0.44 69.64±0.00 58.33±0.34 15
4 LoRA 90.79±0.10 68.71±0.00 82.61±0.00 84.36±0.00 78.41±0.00 76.41±0.00 47.87±0.00 75.76±0.30 90.20±0.00 76.05±0.00 63.60±0.29 65.55±0.32 80.58±0.00 74.77±0.26 65.48±0.42 41.91±0.00 0
AdaLoRA 90.15±0.10 67.35±0.00 81.58±0.29 84.20±0.22 79.55±0.00 76.92±0.00 46.46±0.25 74.71±0.15 90.20±0.00 75.77±0.20 61.76±0.29 63.76±0.00 79.61±0.00 74.59±0.00 65.18±0.00 42.16±0.35 0
BayesTune-LoRA 90.23±0.10 67.12±0.32 80.75±0.00 84.36±0.00 79.55±0.00 76.92±0.00 46.28±0.00 74.71±0.15 90.20±0.00 75.21±0.00 61.35±0.00 63.76±0.00 79.61±0.00 74.59±0.00 65.18±0.00 42.40±0.35 0
FedL2P 90.82±0.11 68.37±0.34 83.54±0.93 85.54±1.19 80.12±0.57 76.41±0.51 48.67±0.80 77.43±0.63 90.20±0.00 76.68±0.21 63.80±1.23 67.78±2.02 81.31±0.73 75.95±0.81 66.51±0.45 43.75±1.84 0
FedP2EFT 90.86±0.36 72.56±0.32 86.96±0.00 91.00±0.39 87.88±0.54 81.37±0.48 52.13±0.00 81.61±0.15 94.77±0.00 79.13±0.20 74.03±0.29 80.32±0.32 83.50±0.00 84.86±0.00 77.08±0.84 68.87±0.35 16
8 LoRA 90.65±0.10 68.71±0.00 83.02±0.29 84.83±0.00 80.68±0.00 77.95±0.00 50.00±0.00 77.43±0.00 90.20±0.00 76.89±0.00 64.42±0.00 66.89±0.32 82.04±0.00 76.58±0.25 66.07±0.00 43.38±0.00 0
AdaLoRA 90.23±0.10 67.35±0.00 80.96±0.29 84.36±0.00 79.55±0.00 76.92±0.00 46.28±0.00 74.71±0.15 90.20±0.00 75.63±0.00 61.76±0.29 63.54±0.32 79.45±0.23 74.59±0.00 65.18±0.00 42.16±0.35 0
BayesTune-LoRA 90.23±0.27 67.35±0.00 81.37±0.51 84.20±0.22 79.55±0.00 76.92±0.00 46.28±0.00 74.61±0.00 90.20±0.00 75.35±0.20 61.55±0.29 63.98±0.32 79.29±0.23 74.59±0.00 65.18±0.00 42.40±0.35 0
FedL2P 90.79±0.10 68.94±0.32 83.85±0.51 85.62±0.81 80.31±1.07 77.43±0.73 49.29±1.25 77.95±0.97 90.42±0.31 76.89±0.00 65.03±1.80 68.46±1.45 81.72±0.46 76.94±1.02 66.67±0.84 45.10±1.93 1
FedP2EFT 90.37±0.36 73.70±0.32 89.23±0.29 93.36±0.39 89.39±0.53 79.66±0.64 51.42±0.67 81.09±0.29 96.08±0.00 78.99±0.34 77.10±0.29 78.52±0.55 82.85±0.23 86.49±0.00 77.38±0.42 67.65±0.00 15
16 LoRA 90.93±0.00 70.75±0.00 85.09±0.00 87.36±0.23 82.95±0.00 80.51±0.00 50.35±0.25 79.83±0.15 90.85±0.00 76.89±0.00 71.78±0.00 71.81±0.95 83.17±0.23 80.18±0.67 67.86±0.00 51.96±0.69 1
AdaLoRA 90.15±0.10 67.58±0.32 80.75±0.00 84.36±0.00 79.55±0.00 76.92±0.00 46.10±0.25 74.71±0.15 90.20±0.00 75.63±0.34 61.55±0.29 63.98±0.32 79.61±0.00 74.59±0.00 65.18±0.00 42.40±0.35 0
BayesTune-LoRA 90.58±0.10 67.35±0.00 81.16±0.29 84.36±0.00 79.55±0.00 76.92±0.00 46.10±0.25 74.61±0.00 90.20±0.00 75.63±0.34 61.35±0.00 63.98±0.32 79.45±0.23 74.59±0.00 65.18±0.00 42.16±0.35 0
FedL2P 91.14±0.30 71.20±0.32 87.17±0.77 87.20±0.67 83.71±0.54 79.32±0.25 51.42±0.67 79.94±0.26 91.72±0.31 77.73±0.91 70.55±2.18 75.17±3.34 82.36±0.23 79.46±0.44 71.43±1.26 55.64±4.81 2
FedP2EFT 88.33±0.70 74.83±0.56 88.41±0.29 93.68±0.23 89.39±0.53 80.51±0.00 51.06±0.44 80.15±0.29 96.51±0.31 80.67±0.00 78.94±1.04 77.63±0.32 82.20±0.23 85.95±0.44 77.38±0.84 68.87±0.69 12
4.1.3 Baselines
Given a base model, we compare FedP2EFT with existing fine-tuning and learning to personalize approaches.
LoRA PEFT. We deploy LoRA (Hu et al., 2022) on all linear layers of the model with a fixed rank $r$.
Non-FL Rank Selection. We compare with AdaLoRA (Zhang et al., 2023) and our proposed LoRA-variant of BayesTune (Kim & Hospedales, 2023), BayesTune-LoRA (Section 3.2), which optimizes $\bm{\lambda}$ separately for each client.
FL to Personalize. We compare with FedL2P (Lee et al., 2023) which trains a MLP federatedly to output per-client learning rates for each LoRA module.

### v2

@@ 3.3 FedP2EFT: FL to Personalize PEFT
Overfitting often occurs when training an effective per-client PS in isolation due to limited client data. Following FedL2P (Lee et al. 2023), we mitigate this by federatedly learning a common PSG that generates client-wise PS. Concretely, we use a one hidden layer multilayer perceptron (MLP) with parameters $\phi$ that takes as input the client meta-data and outputs an estimated PS:
@@
Figure 2: FedP2EFT’s federated training of PSG for each federated round. Details in Section. $3.$
Federated Training of FedP2EFT. Fig. 2 shows the entire FedP2EFT algorithm during federated training. For each federated round, each sampled participating client $i$ receives $\phi$ from the server and loads them into its MLP. They then 1 perform a forward pass of the local train dataset on their base model and a forward pass of the MLP with the resulting features as per Eq. (3.3). 2 The estimated $\bm{\hat{\lambda}}^{i}$ is plugged into our proposed BT-LoRA (Section 3.2) and 3 fine-tuning is performed as per Eq. (1) for $s$ steps (Stage 1). 4 The resulting $\hat{\bm{\lambda}}^{i,s}$ is used as an approximated ground-truth for regression learning of MLP to target the learned $\hat{\bm{\lambda}}^{i,s}$, where $\mathcal{L}_{1}$ is the L1 loss (Stage 2). Finally, 5 $\phi$ is sent back to the server for aggregation. As there is no single aggregation method that outperforms all others in every situation (Matsuda et al. 2022; Chen et al. 2022; Ye et al. 2024), we utilize FedAvg (McMahan et al. 2017). The aggregated $\phi$ is then sent to clients for the next round. To complement Figure 2, we provide Algorithm 1 in Appendix D.
The learned $\phi$ is deployable to any client (seen or unseen) after federated training. Note that unlike FedL2P, which requires federated training for every target rank, FedP2EFT inherits the property of BT-LoRA; federated training is a one-time cost for all ranks $\leq r_{\text{max target}}$.
Inference with FedP2EFT. Fig. 1(b) shows how FedP2EFT personalizes PEFT for each client upon deployment. Given the learned MLP and the client’s base model, 1 the client meta-data are retrieved (Eq. 3.3) and used to generate the client’s PS, $\bm{\lambda}$. 2 Given the client’s resource budget of total rank $r\cdot L$, we take the top-$(r\cdot L)$ largest ranks in $\bm{\lambda}$, freeze them, and initialize our proposed BT-LoRA modules for all layers where $\|\lambda_{l}\|_{1}$>>0. 3 The personalized LoRA ranks are used for fine-tuning before merging back to the base model to obtain the final personalized model. To complement Figure 1(b), we provide Algorithm 2 in Appendix D.
@@ 4 Evaluation
We conduct experiments on multilingual scenarios, where clients with diverse high- and low-resource languages can collaboratively learn how to personalize a given base model to better cater to their language preferences. In all experiments, we divide clients in two pools, seen and unseen, where only the clients in the seen pool actively participate in federated training. We set the maximum number of communication rounds for training the PSG to $150$, randomly sampling $10\%$ of participating clients every round. We use Adam (Kingma and Ba 2015) as the default optimizer for all our experiments. Unless mentioned otherwise, we show results for resource budget $r=2$ where the total rank budget is $r\cdot L$ and leave results for other resource budgets $r=4,8,16$ in the Appendix. In the following subsections, we summarize the FL scenarios that we consider in our experiments, leaving comprehensive details in Appendix A.
4.1 Setup: Tasks, Models, and Datasets
Text Classification. We adopt the pretrained multilingual BERT (Devlin et al. 2018) (mBERT) for all text classification experiments. For datasets, we introduce additional data heterogeneity to the simulated FL setups, XNLI (Conneau et al. 2018) and MasakhaNEWS (Adelani et al. 2023), proposed in PE_FL (Zhao et al. 2024).
For our XNLI setup, we sample 2k instances for train and $500$ for test in each pool. In contrast to PE_FL, which had $15$ clients ($1$ client per language), we divide the data equally among $20$ clients per language. We then adopt the latent Dirichlet allocation (LDA) partition method (Hsu, Qi, and Brown 2019; Yurochkin et al. 2019), $y\sim Dir(\alpha)$, to simulate non-IID label shifts among these clients, with $\alpha=0.5$. Hence, there is a total of $600$ clients ($15$ languages $\cdot$ 2020 clients $\cdot$ 22 pools), consisting of both label and feature heterogeneity.
@@ 4.1 Setup: Tasks, Models, and Datasets
Figure 3: No. of clients by predominant language in our Fed-Aya setup
Instruction-Tuning Generation. We use pretrained MobileLLaMA-1.4B (Chu et al. 2023) and Llama-3.2-3B (Dubey et al. 2024), which are representative of commonly supported model sizes on high-end edge devices (Mehta et al. 2024; Tan et al. 2024; Xu et al. 2024a), and run experiments with them on the Fed-Aya dataset (Ye et al. 2024). Fed-Aya, a real-world FL dataset naturally partitioned by annotator ID, comprises 38 clients, $8$ of whom we designate as our unseen pool. We split each client’s data $80\%/20\%$ for training and testing. Each client has up to $4$ languages and Fig. 3 shows the predominant language (client’s most frequent language) distribution in our setup.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
eng 90.4±0.1 89.9±0.0 89.9±0.0 90.7±0.6 92.0±0.0
som 60.1±0.3 59.4±0.3 59.6±0.3 61.0±1.2 65.5±1.2
run 81.4±0.5 81.2±0.3 81.4±0.0 82.0±0.9 87.8±0.3
fra 88.6±0.0 88.6±0.0 88.6±0.0 89.1±0.7 93.5±0.2
lin 83.5±0.5 82.8±0.0 83.1±0.5 83.9±0.0 88.5±0.9
ibo 79.8±0.2 79.0±0.0 79.0±0.0 79.7±0.2 82.6±0.0
amh 45.7±0.0 45.2±0.0 45.2±0.0 45.7±0.0 52.0±0.2
hau 75.8±0.0 75.0±0.1 75.2±0.0 76.7±1.1 79.7±0.3
pcm 96.0±0.0 96.0±0.0 96.0±0.0 96.0±0.0 97.6±0.3
swa 79.0±0.0 78.6±0.0 78.6±0.0 79.7±0.4 84.7±0.5
orm 64.2±0.0 64.0±0.3 64.0±0.3 64.4±0.3 72.2±1.0
xho 69.1±0.3 69.6±0.0 69.6±0.0 69.1±0.3 76.3±0.6
yor 79.2±0.2 78.9±0.2 79.0±0.0 79.7±0.6 82.1±0.2
sna 78.8±0.0 78.8±0.0 78.8±0.0 78.8±0.0 81.3±0.7
lug 67.6±0.0 67.6±0.0 67.6±0.0 67.9±0.4 69.1±0.4
tir 44.9±0.0 44.9±0.0 44.9±0.0 45.3±0.7 63.5±0.3
Table 1: Mean±SD Accuracy of each language across 3 different seeds for seen clients of our MasakhaNEWS setup ($r=2$, see Appendix Table 7 for other $r$ values). The pretrained model is trained using Standard FL with full fine-tuning and the resulting base model is personalized to each client given a baseline approach.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
eng 90.7±0.0 90.3±0.0 90.3±0.0 90.6±0.1 90.7±0.1
som 68.5±0.3 67.3±0.0 67.3±0.0 68.5±0.3 72.1±0.6
run 82.0±0.5 80.8±0.0 80.8±0.0 82.6±0.0 88.4±0.3
fra 84.4±0.0 84.4±0.0 84.4±0.0 84.4±0.0 88.2±0.4
lin 79.5±0.0 79.5±0.0 79.5±0.0 79.5±0.0 86.4±0.0
ibo 76.8±0.2 76.9±0.0 76.9±0.0 76.4±0.4 81.5±0.4
amh 46.3±0.0 46.3±0.0 45.9±0.2 46.8±0.8 51.1±0.0
hau 75.5±0.1 74.8±0.1 74.6±0.0 76.0±0.6 79.5±0.3
pcm 90.2±0.0 90.2±0.0 90.2±0.0 90.2±0.0 93.7±0.3
swa 75.6±0.3 75.9±0.2 75.5±0.2 75.9±0.4 78.0±0.4
orm 62.0±0.0 61.5±0.3 61.5±0.3 62.2±0.3 73.0±0.5
xho 64.2±0.6 64.2±0.3 63.8±0.0 64.4±0.9 78.5±1.1
yor 80.1±0.0 79.6±0.0 79.5±0.2 80.3±0.2 83.7±0.2
sna 74.6±0.0 74.6±0.0 74.4±0.2 74.8±0.3 80.0±0.4
lug 65.2±0.0 65.2±0.0 65.2±0.0 66.1±0.0 69.6±0.0
tir 41.9±0.0 41.9±0.0 42.6±0.0 41.9±0.0 58.3±0.3
Table 2: Mean±SD Accuracy of each language for unseen clients of our MasakhaNEWS setup ($r=2$, see Appendix Table 8 for other $r$ values). The pretrained model is trained using Standard FL with full fine-tuning and the resulting base model is personalized to each client given a baseline approach.
4.2 Setup: Complementary Approaches
FedP2EFT is compatible with off-the-shelf models and models trained with existing FL approaches. Concretely, given a pretrained model, we obtain a base model using one of the following approaches:
Standard FL. We further train the pretrained model federatedly on the seen pool, either using existing PEFT methods or full fine-tuning (Ye et al. 2024; Sun et al. 2024),
Personalized FL (pFL). We adopt two recent pFL works: i) FedDPA-T (Yang et al. 2024b), which learns per-client personalized LoRA modules in addition to global LoRA modules, and ii) DEPT (SPEC) (Iacob et al. 2025), which learns per-client personalized token and positional embeddings while keeping the rest of the model shared. The base model hence differs for each client.
@@ 4.2 Setup: Complementary Approaches
4.3 Setup: Baselines